Consider a binomial distribution with 10 trials What is the expected value of this distribution if the probability of success on a single trial is 0.5?
What is the chance of rolling a 6 4 times in a row?
The chance of rolling it each time is (1/6). So the chance of rolling it 4 times in a row is
(1/6) x (1/6) x (1/6) x (1/6) = (1/1296) = 0.0007716 = 0.0772 % (rounded)
The chance of rolling it each time is the same. It's not affected by what came before,
since dice have no memory.
Given that p equals log2q find an expression in terms of p for log2 routeq and log2 8q?
log2 sqrt(q) = p/2
log2 8q = 3 + p
And it is root, not route.
The distribution of sample means will not be normal if the number of samples does not reach 30.
If you roll two dice what is the probability of getting an odd number?
The probability of getting an odd number when you roll one die is 1 in 2. In order to get an odd sum with two dice, one of them has to be odd and one of them has to be even. The probability of rolling an odd sum is still 1 in 2, since each die is unrelated in probability to the other.
What is the probablity of rolling the sum of 12 on 2 dice?
There are 36 outcomes for rolling 2 dice, and there is 1 way that a 12 can occur which is 6,6. So, the probability of rolling the sum of 12 on 2 dice is 1/36.
What is the probability of rolling the sum of 12 with 2 dice?
Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.
Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.
Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.
Since you need a specific number - the number 6 - twice, the probability is (1/6)2, which is equal to 1/36.
What is the difference between z-score and t-score?
T-scores and z-scores measure the deviation from normal. The normal for T-score is 50 with standard deviation of 10. if the score on t-score is more than 50, it means that the person scored above normal (average), and vise versa. The normal for Z-score is 0. If Z-score is above 0, then it means that person scored above normal (average), and vise versa.
What is the purpose of snowball sampling?
Snowball sampling can be used in a number of circumstances.
Sometimes the researcher may be able to identify only a small number of individuals in the target group but these people are likely to known others in the target group, and people that they know will also know more such people-and so on. By including those few initial people in her sample, and asking them to recruit others, the researcher may be able to achieve a useable sample size.
Then, sometimes, a researcher is actually more interested in the connections between people, or in other words, social networks. In this case, snowball sampling can be used to identify those connections.
Please see the link.
What is the probability of drawing 2 hearts and then a spade in that order from a deck of 52 cards?
In order to determine the probability of drawing 2 hearts and then a spade, in that order, from a deck of 52 cards, start by considering the first card.
The probability of drawing a heart is 1 in 4.
Since you have now reduced the number of hearts and the number of cards in the deck by one, the probability of drawing another heart is 4 in 17.
Since you have further reduced the number of cards by one, the probability of drawing a spade is 13 in 50.
Multiply these probabilities together, (1/4) (4/17) (13/50), and you get about 0.0153, or about 153 in 10000.
What is the probability of rolling an odd number and a prime number?
With a standard 6 faced die, numbered 1 through 6, there are three odd numbers, 1, 3, and 5, of which 3 and 5 are prime, and there is one even prime, 2. The probability, then, of rolling an odd number which is prime is 1 in 3, i.e. 3 or 5 out of a possible result set of 1, 2, 3, 4, 5, and 6.
If you flip a coin 4 times how many possible outcomes are there?
There are 24 = 16 ordered outcomes, that is outcomes in which the order of the results is relevant. If not, there are 5 outcomes (0 heads, 1 head, 2 heads, 3 heads and 4 heads).
How many ways are there to select 10 balls from unlimited nunber of red white blue and green balls?
in those 10 balls we can have any combination...I am using here multinomial...
X=No. of red balls(0,1...10)
y=No. of green balls(0,1...10)
z=No. of white balls(0,1...10)
we need co-efficient of x^10 in (x^0+x^1+...+x^10)^3
that is (10+3-1)C3=12C3
Arrangement is also considered here
What is the probability of not tossing three heads with three fair coins?
The probability of something NOT happening is the complement of the probability of something happening. Since the probability that you DO have 3 heads is 1/8 (that is, 1/2 cubed), the complement is 1 - 1/8 = 7/8.
What is the Probability of drawing two spades out of 52 cards?
The probability of drawing a spade in a standard 52 card deck is 13 in 52, or 1 in 4. The probability of drawing a second spade, assuming the first spade was not replaced back into the deck, is 12 in 51. The probability, then, of drawing two spades is the product of those two probabilities, or 12 in 204, or 1 in 17.